OpenStudy (anonymous):

Determine if (1/2,-4) is a solution to 4x-3y=-10.Please show all your work.

OpenStudy (anonymous):

graph line, plot point. point on line? plug (x,y) into (x,y) of equation. does left hand side = right hand side? 4(1/2) - 3(-4) = -10, for instance will -10= -10?

OpenStudy (anonymous):

So is it a solution?

OpenStudy (anonymous):

No, 4(1/2) - 3(-4) = 2+12 = 14. 14 is not -1, so (1/2, -4) is not a solution.

OpenStudy (anonymous):

is not -10*